A solution of natural water contains 1.00 x 10-3 M Mg2+. Calculate the value of pMg2+ after the following additions of 0.00200 M EDTA to an initial volume of 10.00 mL of the Mg2+ solution. A borax buffer is used to maintain the pH at 10.70 for all the titration.

a) 0.00 mL b) 2.00 mL c) 3.50 mL d) 5.00 mL e) 7.50 mL f) 10.00 mL
Answer #1

The magnesium-EDTA titration is a complexometric titration where Mg2+ forms a complex with EDTA. EDTA is a multidentate (hexadentate ligand). Therefore the complex formation takes place in steps. Therefore we need to consider the conditional formation constant for the complex formation.

If Kf is the formation constant for the metal EDTA complex given by

MgY-

But we need to consider the conditional formation constants \alphaY4- and \alphaMg2+

\alphaY4- = \frac{[Y^{4-}]}{C_{EDTA}}

\alphaMg2+ = \frac{[Mg^{2+}]}{C_{Mg^{2+}}}

Therefore, we get Kf" = Kf\ast\alphaY4-\ast\alphaMg2+.

For pH 10 \alphaY4- = 3.8*10(-8) and \alphaMg2+ = 0.984

Equivalence volume is given by V=(001*1 0.002 =5mL

i) For 0 mL of EDTA added,

[Mg2+] is 0.001.

\RightarrowpMg = -log([Mg2+ ])

[Mg2+ ] = \alphaMg2+ * CMg2+

            = 0.984 * 0.001= 0.0984

pMg= 3.007

ii)For 2mL

Amount of Mg2+ consumed = amount of EDTA added

Amount of Mg2+ left = (0.001*10-0.002-2) =0.0005 M

Using \alphaMg2+ , [Mg2+] = \alphaMg2+ * CMg2+ = 0.0005*0.984= 0.000492

pMg = -log([Mg2+ ]) = 3.30

iii)For 3.5mL

Amount of Mg2+ left = (0.001-10-0.002+3.5) = 0.00022

[Mg2+ ]= \alphaMg2+ * CMg2+ = 0.984 * 0.00022 = 0.0002186

pMg = 3.66

iv)For 5 mL

Here 5 mL is the equivalence volume.

Therefore,

amount of Mg2+ left =( Initial moles of Mg2+) ÷(Total Volume)

                            = (0.001*10)/(15) =0.00066

[Mg2+ ]= 0.00066*0.984 = 0.000656

pMg = 3.18